Math 042: Exam I Date: Feb 24, 2010

Name: Section:

1. Simplify: 8a^{3} − 3a^{2} − (4a^{2} − 5) Your answer: .

2. Solve for x: −3x + 5 = −5x − 8 Your answer: .

3. Solve for x: \frac{4}{3}x − 6 = \frac{3}{2}x Your answer: .

4. Simplify: \frac{(6x^{2}y/11)}{(9y^{2}/22)} Your answer: .

5. Evaluate: −a^{2} + 3b^{3} for a = 5, b = −3

(a) 106 (b) -106 (c) 56 (d) -56 6. Simplify: \frac{(3xy)^{2}}{9x^{2}y^{3}} Your answer: .

7. Solve the inequality: −3(y − 1) < 12 Your answer: .

8. Solve for c: f = \frac{9}{5}c + 32

(a) \frac{5f−32}{9} (b) \frac{5(f−32)}{9} (c) \frac{5f}{9} − 32 (d) \frac{9}{5}f − 32 9. Solve for x: ax + d = cx − b Your answer: .

10. 120% of what number is 30? Your answer: .

11. 24 is what percent of 80? Your answer: .

12. 6 times the sum of a number and its cube can be represented by

(a) 6n + n^{2} (b) 6(n + n)^{3} (c) 6 + n + n^{3} (d) 6(n + n^{3})

13. The length of a rectangle is 5 less than twice its width. What can be used to represent its perimeter?

Your answer: .

14. An RCA employee earned $280 a week. If $52 is deducted for taxes, $35 for health benefits and $25 for pension, what percent of the salary is left?

(a) 168% (b) 112% (c) 60% (d) 40%

15. If $1300 is invested at t% and $600 is invested at (t − 3)% what can be used to represent the interest drawn in one year?

Your answer: .